[figure: 9.7 cm]

ELECTRON PARAMAGNETIC RESONANCE OF Gd^{3+} IN TlGdSe_{2}

M. Duczmal^{a}, E. Mosiniewicz-Szablewska^{b} and S. Pokrzywnicki^{c} aWrocław University of Technology, Department of Chemistry, Wyb. Wyspiańskiego 27, 50-370 Wrocław, Poland

bInstitute of Physics, Polish Academy of Science, Al. Lotników 32-46, 02-668 Warszawa

cTechnical University of Opole, Institute of Mathematics, Physics and Chemistry, ul. Luboszycka 7, 45-036 Opole, Poland

TlGdSe_{2} crystallizes in the rhombohedral α-NaFeO_{2} type structure (R¯3m). It is a NaCl-related ordered structure with the thallium and gadolinium layers separated by the anion layers. The electron paramagnetic resonance shows one asymmetric line with the temperature dependent width. The linewidth remains nearly constant when the temperature is lowered from room temperature and steep narrowing is observed below 10 K. The half-power half-width calculated assuming the purely dipolar interactions (212 mT) is bigger than the experimental one (155 mT at room temperature). The difference was ascribed to the short range exchange interactions.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Marek Duczmal

Address for correspondence : Politechnika Wrocławska Instytut Chemii Nieorganicznej i Metalurgii Pierwiastków Rzadkich Wyb. Wyspiańskiego 27 50-370 Wrocław

Email address : duczmal@ichn.ch.pwr.wroc.pl